OpenStudy (anonymous):

show that each number is rational by writing it in the form a/b, where a and b are intergers. 1. 0.37 2. 2.1

OpenStudy (anonymous):

plz help

OpenStudy (zzr0ck3r):

\(a.bcdef...=\frac{a}{1}+\frac{b}{10}+\frac{c}{100}+\frac{d}{1000}+\frac{e}{10000}+\frac{f}{100000}...\)

OpenStudy (zzr0ck3r):

example: \(1.34=1+\frac{3}{10}+\frac{4}{100}\\=\frac{100}{100}+\frac{3*10}{100}+\frac{4}{100}=\frac{100+30+4}{100}\\=\frac{134}{100}\\=\frac{67}{50}\)

jimthompson5910 (jim_thompson5910):

`0.37` means `37 hundredths` since that 7 is in the hundredths place so you write `37` over 100 to get \[\Large \frac{37}{100}\] this means \[\Large \frac{37}{100} = 0.37\] I'll let you do the other one